A farmer is going to divide her 60 acre farm between two crops. Seed for crop A costs $20 per acre. Seed for crop B costs $40 per acre. The farmer can spend at most $2200 on seed.
If crop B brings in a profit of $190 per acre, and crop A brings in a profit of $110 per acre, how many acres of each crop should the farmer plant to maximize her profit?

Answer by ikleyn(52779)

.
Let X = acres for A  and  Y = acres for B.


The farmer wants to maximize his profit  

    P(X,Y) = 110*X + 190*Y


under these restrictions

    X + Y = 60,

    20*X + 40*Y <= 2200,

    X >= 0,  Y >= 0.



The feasible domain is shown in the Figure below.






Plot X + Y = 60 (red)  and  20X + 40Y = 2200 (green)



The feasibility domain is the quadrilateral in QI below the red and green lines.


Corner points of the quadrilateral (its vertices) are

    P1 = (0,0)

    P2 = (0,55)   (y-intercept of the green line)

    P3 = (10,50)  (intersection point of red and green line).

    P4 = (60,0)    (x-intercept of the red line)


The values of the profit function in the corner points are:


    P1:  P( 0, 0) = 0

    P2:  P( 0,55) = 110*0  + 190*55 = 10450

    P3:  P(10,50) = 110*10 + 190*50 = 10600

    P4:  P(60,0)  = 110*60 + 190*0  = 6600.


The maximal value of the profit function is achieved at the points P3.

It means that the optimal solution is 10 acres for A and 50 acres for B.


It gives the profit of 16000 dollars.

Solved.

--------------------

To see other similar problems solved by the Linear Programming method,  look into the lesson
    - Solving minimax problems by the Linear Programming method
in this site.

Also,  look into the solutions of other similar problems in the archive to this forum under the links

https://www.algebra.com/algebra/homework/Coordinate-system/Coordinate-system.faq.question.1134444.html
https://www.algebra.com/algebra/homework/Graphs/Graphs.faq.question.1131906.html
https://www.algebra.com/algebra/homework/coordinate/word/Linear_Equations_And_Systems_Word_Problems.faq.question.1131043.html
https://www.algebra.com/algebra/homework/word/finance/Money_Word_Problems.faq.question.1129285.html
https://www.algebra.com/algebra/homework/Finance/Finance.faq.question.1128383.html
https://www.algebra.com/algebra/homework/Linear-equations/Linear-equations.faq.question.1123217.html
https://www.algebra.com/algebra/homework/playground/test.faq.question.1112482.html
https://www.algebra.com/algebra/homework/Finance/Finance.faq.question.1102103.html